OpenStudy (anonymous):

how do i get -100000 into polar form to find all 5 complex 5th roots?

OpenStudy (anonymous):

(-10)^5=-100000

OpenStudy (anonymous):

yeah i know but i need to show all the work using demoivres theorem

OpenStudy (anonymous):

so i need to get it into polar form....like cos sin

OpenStudy (anonymous):

\[x^5=-100000\] so that's the question right?

OpenStudy (anonymous):

and now you want to use laws of demoivre to get all the answer in form of polar coordinates?

OpenStudy (anonymous):

Find all the COMPLEX FIFTH ROOTS OF -100000. Write the roots in rectangular form.

OpenStudy (anonymous):

well so I guess my equation above is correct, excuse me, my english is not flawless but that's what I would do, because the equation above has 5 roots in the complex plane. So you can tell already from the equation above, that this can be written as: \[x^5=-100000 +0i \] which means it is located on the negative real axis, therefore an argument of this can be found, an argument is \(\pi\)

OpenStudy (anonymous):

im looking for 5 roots...five different answers.

OpenStudy (anonymous):

\[ \phi = \frac{\pi}{5} + \frac{2\pi k }{5} \] k \( \in \mathbb{N}\)

OpenStudy (anonymous):

gives you five answer, choose k=0, k=1, k=2, K=3, k=4

OpenStudy (anonymous):

what is k

OpenStudy (anonymous):

i have no idea where that comes from im just going to have to do this myself. thanks anyways.

OpenStudy (anonymous):

a number, you can choose it for yourself, just like in trigonometry when you are looking for all the various angles that satisfy an equation.

OpenStudy (anonymous):

well it's the legit way, what you are looking for is the following \[ \sqrt[5]{100000}( \text{cis} \phi) \]

OpenStudy (anonymous):

this is the polar coordinate form, the result can be both, either polar or rectangular.

OpenStudy (anonymous):

\[ 10 ( \cos \phi + i\sin \phi) \]

OpenStudy (anonymous):

ive got it.

OpenStudy (anonymous):

good (-: